OpenStudy (anonymous):

Find the common difference of the arithemetic a1= -22 and a12 =77

OpenStudy (anonymous):

@thomaster

OpenStudy (***[isuru]***):

Hi, do u know the formula to find the nth term of an arithmetic progression ?

OpenStudy (***[isuru]***):

\[T _{n} = a + (n-1)d\] ?

OpenStudy (anonymous):

TEA THATS IT

OpenStudy (anonymous):

no wait no

OpenStudy (***[isuru]***):

wt do y mean by that ? Do u know a formula which is different to this one ?

OpenStudy (***[isuru]***):

anyway... ya can solve this easily with the one i mention above u can find the n th term of an arithmetic progression by ...\[T _{n} = a + (n-1)d\] where a is the first term... d is the common difference... n / Tn is the n th term now in ur question a1 is given as -22 so ... we know the first term of this progression. and the value of 12 th term is also given as 77 now.. if we use that formula to find the 12th term... \[T _{12} = a + (12 -1)d\] we know the value of 12 th term and the first term.... so we can rewrite this as...\[77 = -22 + (12 -1)d\]\[99 = 11d\]\[d = 9\] hope this will help ya out!!!

OpenStudy (anonymous):

Thanks so much

OpenStudy (***[isuru]***):

u r welcome!!